# Continuum models describing probabilistic motion of tagged agents in exclusion processes

### v1

@@ I Introduction
Lattice-based interacting random walk models, including exclusion processes and generalisations thereof, are often used to study collective motion in populations of biological cells [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Similar lattice-based and lattice-free stochastic models that incorporate crowding mechanisms have been used to study animal and plant dispersal in the context of spatial ecology [16, 17, 18, 19, 20]. Lattice-based exclusion process models enforce each lattice site to be occupied by, at most, a single agent to capture experimentally observed crowding effects [21, 22]. Potential motility or proliferation events that would place more than one agent at the same site are aborted. These models can be used to mimic data generated during cell biology experiments by generating simulation-based snapshots and movies that are directly comparable with experimental images [23]. These simulation-based outputs are characterised by fluctuations and stochasticity that are also present in experimental data. Beyond simply generating simulation-based images and movies, interacting random walk models can be used to generate ensemble data by considering a suite of identically prepared realisations and averaging over these stochastic realisations to give an averaged, population-level description [22]. This kind of ensemble data can also be modelled using partial differential equation (PDE)-based descriptions for the average population density $C(x,y,t)\in[0,1]$ that are obtained by applying the mean-field approximation [24].
Experimental observations often encompass data across multiple scales. Populations of simulated individuals undergoing motility and carrying capacity-limited proliferation typically lead to moving population fronts, as illustrated by the schematic front in Figure 1(a) that moves in the positive $x$-direction as the population invades adjacent regions. Both in vivo and in vitro experimental observations describe this kind of population front movement as illustrated in Figure 1(b)-(c) [25, 26]. It is also possible to use fluorescent labels to highlight the motion of individual cells within the invading populations, such as the trajectories superimposed in Figure 1(b)-(c) [25, 26]. In these two cases, we see that cells close to the low-density leading edge of the invading population undergo random motion that is biased in the same direction as population-level front motion. These labelled cells at the low-density leading edge undergo longer trajectories than cells in higher-density regions behind the leading edge.
@@
Figure 1: Schematic and experimental motivation. (a) Schematic scaled density profile showing the spatial expansion of a population of cells, undergoing migration and proliferation, leading to the macroscopic propagation of a density front in the positive $x$-direction. (b) In vivo tagged cell trajectories reported by Druckenbrod and Epstein [25]. The direction of the population density front motion is shown with the blue arrow and individual cell trajectories within the population are given by the red and green traces. (c) In vitro tagged cell trajectories in a wound healing experiment reported by Cai et al. [26]. The direction of the population density front motion is shown with the blue arrow and individual trajectories confirm that cells at the edge of population front are biased to move in the same direction as the population front. All images are reproduced with permission.
@@
In the present study we take a different approach and derive a different, more informative macroscopic PDE that gives a probabilistic prediction of the motion of tagged agents. We derive new PDE models describing the evolution of the probability density function for tagged agents $P^{(s)}(x,y,t)\geq 0$. The general framework applies to a population of agents composed of $S\in\mathbb{Z}^{+}$ distinct subpopulations of agents on the lattice, where each subpopulation undergoes biased motility and proliferation at potentially distinct rates. The system of PDEs for $P^{(s)}(x,y,t)\geq 0$ for $s=1,2,\ldots,S$ can be written as a system of conservation equations with a natural physical interpretation that we will explore. We obtain numerical solutions of the system of PDEs, giving estimates of $P^{(s)}(x,y,t)\geq 0$ for $s=1,2,\ldots,S$, and show that these solutions accurately describe stochastic simulation data, including both the average trends in the trajectory data, as well as capturing the observed variability in the simulation data.
II Methods
II.1 Stochastic model and continuum-limit description: Single species $S=1$
@@
We consider a stochastic agent-based model (ABM) representing an asymmetric exclusion process, simulated using a random sequential update method [28] on a two-dimensional regular lattice with spacing $\Delta$. In each time step, of duration $\tau$, all agents have the opportunity to move with probability $M$. A motile agent at $(x,y)$ steps to $(x,y\pm\Delta)$ with probability $(1\pm\rho_{y})/4$, or to $(x\pm\Delta,y)$ with probability $(1\pm\rho_{x})/4$, where $\left|\rho_{x}\right|\leq 1$ and $\left|\rho_{y}\right|\leq 1$. Here $\rho_{x}$ and $\rho_{y}$ are constant bias parameters that control the degree of motility bias, and we note that setting $\rho_{x}=\rho_{y}=0$ leads to unbiased motility. In each time step, of duration $\tau$, all agents have the opportunity to proliferate with probability $Q$. A proliferative agent at $(x,y)$ attempts to place a daughter agent at sites $(x\pm\Delta,y)$ or to $(x,y\pm\Delta)$. Each of the four possible target sites are chosen with equal probability $1/4$. Any potential motility or proliferation event that would place an agent on an occupied site is aborted.
@@
II.2 Stochastic model and continuum-limit description: Multiple species $S>1$
@@
II.3 Tagged agents
We now describe how to obtain a macroscopic model describing the motion of individual agents within the population using a novel probabilistic framework. To proceed, we suppose that one of the individual agents in the population is tagged at time $t=0$ and its location tracked through time. Let $P(x,y,t)$ denote the probability density function for the agent’s location at time $t$. We now derive a continuum-limit PDE for $P(x,y,t)$ via a similar method to that used to derive the PDE for $C(x,y,t)$.
@@ II.3 Tagged agents
where ${T(x,y,t)=\sum_{s=1}^{S}C^{(s)}(x,y,t)}$ is the total density. The three terms in these expressions for the multi-species flux have a similar physical interpretation as for the single species case described above, and when $S=1$ the multi-species PDE model reduces to the single species case given by Eq. (8).
@@
It is worth noting the parallels between the flux terms for the probability density function for tagged agent location in Eq. (II.3) and those for macroscopic agent density in Eq. (II.2). Effectively, a tagged agent can be interpreted as a separate species, moving within a population with total macroscopic density $T(x,y,t)$. However, there are important differences between the variables $C^{(s)}$ and $P^{(s)}$. Since we have an exclusion process, a key property of the ABM is that the occupancy of individual sites cannot exceed $1$, and in the continuum limit this means that we have $C^{(s)}(x,y,t)\in[0,1]$ for $s=1,2,\ldots,S$, with the additional constraint that ${T(x,t)=\sum_{s=1}^{S}C^{(s)}(x,y,t)}\in[0,1]$. These properties do not hold for the probability density functions for tagged agent locations since we have $P^{(s)}(x,y,t)\geq 0$ with the constraint that ${\int_{\Omega}P^{(s)}(x,y,t)\,\textrm{d}x\,\textrm{d}y}=1$, for $s=1,2,\ldots,S$. There is no upper bound on $P^{(s)}$ and indeed to model a tagged agent initially located at $(x,y)=(x_{0},y_{0})$, we use an initial condition $P^{(s)}(x,y,0)=\delta(x_{0},y_{0})$ where $\delta(\cdot)$ is the Dirac delta function.
II.4 Numerical methods

### v2

@@ 1 Introduction
Lattice-based interacting random walk models, including exclusion processes and generalisations thereof, are often used to study collective motion in populations of biological cells [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Similar lattice-based and lattice-free stochastic models that incorporate crowding mechanisms have been used to study animal and plant dispersal in the context of spatial ecology [16, 17, 18, 19, 20]. Lattice-based exclusion process models enforce each lattice site to be occupied by, at most, a single agent to capture empirically observed crowding effects [21, 22]. Potential motility or proliferation events that would place more than one agent at the same site are aborted. These models can be used to mimic data generated during cell biology experiments by generating simulation-based snapshots and movies that are directly comparable with experimental images [23]. These simulation-based outputs are characterised by fluctuations and stochasticity that are also present in experimental data. Beyond simply generating simulation-based images and movies, interacting random walk models can be used to generate ensemble data by considering a suite of identically prepared realisations and averaging over these stochastic realisations to give an averaged, population-level description [22]. This kind of ensemble data can also be modelled using partial differential equation (PDE)-based descriptions for the average population density that are obtained by applying the mean-field approximation [24].
Experimental observations often encompass data across multiple scales. Populations of simulated individuals undergoing motility and carrying capacity-limited proliferation typically lead to moving population fronts, as illustrated by the schematic front in Figure 1(a) that moves in the positive $x$-direction as the population invades adjacent regions. Both in vivo and in vitro experimental observations exhibit this kind of population front movement as illustrated in Figure 1(b)-(c) [25, 26]. For example, Figure 1(b) shows an experimental image of neural crest cells (NCC) cells moving within live, intact tissues during the development of the enteric nervous system (ENS). Individual NCCs are motile and proliferative, leading to a moving front as the population invades adjacent gut tissue [25]. This experiment involves tagging a small subset of NCCs with fluorescent labels within four different regions of the three-dimensional tissue (Regions I–IV in Figure 1(b)). Region I is at the leading edge of the invading population, whereas Region IV is furthest behind the leading edge. Superimposing each recorded trajectory within the particular region, shown in the insets labelled I–IV, shows that cells in region I undergo random biased motion where the trajectories are directed in the same direction as the macroscopic wavefront motion. In contrast, cells in Regions II, III and IV appear to follow random, undirected trajectories. While it is well understood that NCCs are both motile and proliferative [25], the precise details of the motility mechanism remain relatively poorly understood. The experimental image in Figure 1(c) shows trajectories obtained from a scratch assay, where motion of fibroblast cells is constrained to a two-dimensional plastic substrate [26]. Here, in the simpler two-dimensional geometry, the motion of individual cells can be tracked more easily, giving rise to the trajectories shown in this image. In this experiment, fibroblast cells undergo combined migration and proliferation, leading to the macroscopic expansion of the population, and the trajectories show that cells towards the leading edge are biased to move in the same direction as the front. While the experimental images and data contained within Figure 1 highlight the importance of labelling and tracking, these data alone does not provide quantitative insight into the underlying mechanisms that drive these experimental observation. Here we develop quantitative mathematical and computational tools that are motivated by these experimental observations with the aim of developing new modelling tools that can be used to provide insight into these types of experiment.
@@
Figure 1: Schematic and experimental motivation. (a) Schematic scaled density profile showing the spatial expansion of a population of cells, undergoing migration and proliferation, leading to the macroscopic propagation of a density front in the positive $x$-direction. (b) In vivo tagged cell trajectories reported by Druckenbrod and Epstein [25]. The direction of the population density front motion is shown with the blue arrow and individual cell trajectories within the population are given by the red and green traces. (c) In vitro tagged cell trajectories in a wound healing experiment reported by Cai et al. [26]. The direction of the population density front motion is shown with the blue arrow and individual trajectories show that cells at the edge of population front are biased to move in the same direction as the population front. All images are reproduced with permission.
@@
In the present study we take a different approach and derive a different, more informative macroscopic PDE that describes the evolution of the probability density function (PDF) for tagged agents $P^{(s)}(x,y,t)\geq 0$, giving a probabilistic prediction of their movement. The general framework applies to a population of agents composed of $S\in\mathbb{Z}^{+}$ distinct subpopulations of agents on the lattice, where each subpopulation undergoes biased motility and proliferation at potentially distinct rates. The system of PDEs for $P^{(s)}(x,y,t)\geq 0$ for $s=1,2,\ldots,S$ can be written as a system of conservation equations with a natural physical interpretation that we will explore. We obtain numerical solutions of the system of PDEs, giving estimates of $P^{(s)}(x,y,t)\geq 0$ for $s=1,2,\ldots,S$, and show that these solutions accurately describe stochastic simulation data, including both the average trends and observed variability in the trajectory data. Our study focuses on comparing synthetic data from the stochastic model with solutions of the PDE. While strongly motivated by experimental conditions such as those shown in Figure 1, the derivation of a tractable PDE model and validation against synthetic data is an important contribution in itself, providing an analytical framework for applying to experimental trajectory data.
Our approach has some parallels with methods, such as Kriging, Brownian bridges and kernel density estimators, used in animal movement ecology to infer trajectories or home-range estimates from telemetry data [28, 29]. However, our focus is different in that we specifically account for interactions between multiple individuals in a population, rather than assume they are moving independently, and for proliferation of individuals.
2 Model
In this section, we define a stochastic, agent-based model (ABM) for a population of cells undergoing movement and proliferation in a two-dimensional space. This could represent a specific experimental setup, such as that shown in Figure 1b-c. We assume that the cells move on a regular lattice and that crowding effects mean that no more than one cell can occupy a single lattice site. The population may consistent of one or more than subpopulation, where each subpopulation is characterised by four parameters: motility rate, proliferation rate, horizontal direction bias, and vertical directional bias.
From the ABM, we derive an approximate continuum description that takes the form of a set of reaction-advection-diffusion PDEs for mean agent density. The PDEs are nonlinear due to the interactions between agents. The ABM and corresponding PDEs for agent density have been studied previously [30, 27]. The novel contribution of this work is to derive a corresponding PDE for the PDF for the locations of a subset of tagged agents with the population(s).
2.1 Stochastic model and continuum-limit description: Single species $S=1$
@@
Consider a stochastic ABM representing an asymmetric exclusion process, simulated using a random sequential update method [31] on a two-dimensional regular lattice with spacing $\Delta$. In each time step, of duration $\tau$, all agents have the opportunity to move with probability $M$. A motile agent at $(x,y)$ steps to $(x,y\pm\Delta)$ with probability $(1\pm\rho_{y})/4$, or to $(x\pm\Delta,y)$ with probability $(1\pm\rho_{x})/4$, where $\left|\rho_{x}\right|\leq 1$ and $\left|\rho_{y}\right|\leq 1$. Here $\rho_{x}$ and $\rho_{y}$ are constant bias parameters that control the degree of motility bias, and we note that setting $\rho_{x}=\rho_{y}=0$ leads to unbiased motility. In each time step, of duration $\tau$, all agents have the opportunity to proliferate with probability $Q$. A proliferative agent at $(x,y)$ attempts to place a daughter agent at sites $(x\pm\Delta,y)$ or to $(x,y\pm\Delta)$. Each of the four possible target sites are chosen with equal probability $1/4$. Any potential motility or proliferation event that would place an agent on an occupied site is aborted.
@@
2.2 Stochastic model and continuum-limit description: Multiple species $S>1$
@@
Note that, in general, Eqs. (1) and (4) are nonlinear PDEs as there are nonlinear terms in the expressions for the flux $\bf J$} and proliferation rate $\mathcal{S}$. These arise due to the interactions between agents representing crowding effects. In the special case where there is no directional bias and no proliferation (i.e. $v_{x}=v_{y}=\lambda=0$), Eq. (1) for the single-species cases reduces to the linear diffusion equation, but Eq. (4) for the multi-species cases is still nonlinear [30].
2.3 Tagged agents
We now describe how to obtain a macroscopic model describing the motion of individual agents within the population using a novel probabilistic framework. To proceed, we suppose that one of the individual agents in the population is tagged at time $t=0$ and its location tracked through time. Let $P(x,y,t)$ denote the PDF for the agent’s location at time $t$. We now derive a continuum-limit PDE for $P(x,y,t)$ via a similar method to that used to derive the PDE for $C(x,y,t)$.
@@ 2.3 Tagged agents
where ${T(x,y,t)=\sum_{s=1}^{S}C^{(s)}(x,y,t)}$ is the total density. The three terms in these expressions for the multi-species flux have a similar physical interpretation as for the single species case described above, and when $S=1$ the multi-species PDE model reduces to the single species case given by Eq. (8). Like the PDEs for agent density $C$, Eqs. (8) and (10) are both nonlinear PDEs due to the nonlinearities in the flux $\bf J$}.
@@
It is worth noting the parallels between the flux terms for the PDF for tagged agent location in Eq. (2.3) and those for macroscopic agent density in Eq. (2.2). Effectively, a tagged agent can be interpreted as a separate species, moving within a population with total macroscopic density $T(x,y,t)$. However, there are important differences between the variables $C^{(s)}$ and $P^{(s)}$. Since we have an exclusion process, a key property of the ABM is that the occupancy of individual sites cannot exceed $1$, and in the continuum limit this means that we have $C^{(s)}(x,y,t)\in[0,1]$ for $s=1,2,\ldots,S$, with the additional constraint that ${T(x,t)=\sum_{s=1}^{S}C^{(s)}(x,y,t)}\in[0,1]$. These properties do not hold for the PDFs for tagged agent locations since we have $P^{(s)}(x,y,t)\geq 0$ with the constraint that ${\int_{\Omega}P^{(s)}(x,y,t)\,\textrm{d}x\,\textrm{d}y}=1$, for $s=1,2,\ldots,S$. There is no upper bound on $P^{(s)}$ and indeed to model a tagged agent initially located at $(x,y)=(x_{0},y_{0})$, we use an initial condition $P^{(s)}(x,y,0)=\delta(x_{0},y_{0})$ where $\delta(\cdot)$ is the Dirac delta function.
For details of the methods used to simulate the ABM and numerically solve the PDE model, see Supplementary Material [34]. Matlab software to reproduce the results in this study is publicly available at [35]. We encourage readers can use this software directly to replicate the results presented in this study, or to adapt the software and explore different scenarios, such as working with different initial conditions, parameter values, simulation durations.
3 Results
In section 2, we presented the model derivation for the general multi-species case ($S>1$) in two spatial dimensions in order to provide a comprehensive theory. In this section, we provide numerical comparisons of stochastic ABM simulations and solutions of the PDE model, focusing on the single-species case ($S=1$) and the situation where lattice occupancy is independent of the $y$ coordinate (see Supplementary Material [34] for details). We therefore drop the superscript $(s)$ and use $C(x,t)$ and $P(x,t)$ respectively to denote average agent density and PDF for tagged agents at location $x$ and time $t$. We systematically explore the behaviour of the models with and without direction bias, and with and without agent proliferation. Our results illustrate the qualitative types of behaviour that can occur in different situations and test the accuracy of the PDE model approximation $P(x,t)$ for the distribution of tagged agent locations in stochastic simulations.
3.1 Unbiased motility and no proliferation
3.2 Biased motility and no proliferation
When there is a directional bias in movement in the positive $x$ direction with $\rho_{x}>0$, the macroscopic agent density becomes skewed to the right (Figure 2d), as is well known [27]. Again, we see that averaged simulation data for the density matches the solution of the mean-field density, including the skew in the density profile. The distribution of tagged agents initially near the right-most leading edge drifts in the positive $x$-direction due to the motility bias, and diffuses over time (Figure 2e, yellow). Tagged agents initially in the centre $x=0$ and at the left-most leading edge drift more slowly and do not spread out as rapidly due to crowding effects (Figure 2e, red, blue). Again, the PDE gives a good approximation to the distribution of tagged agent locations in the ABM (Figure 2e-f).
Figure 2: Comparison of discrete and continuum models for a motile population: (a)-(c) without bias or proliferation; (d)-(f) with bias and without proliferation; (g)-(i) without bias and with proliferation; (j)-(l) with bias and proliferation. Left column of plots show agent density $C(x,t)$ at $t=300$; middle column shows distribution of the location at $t=300$ of tagged agents initially located near the left-hand leading edge ($x_{0}=-18$, blue), in the centre of the population ($x_{0}=0$, red) and near the right-hand leading edge ($x_{0}=18$, yellow); right column shows the median and 90% PrI of tagged agent locations over time time according to the ABM (thick solid curve = median, shaded band = 90% PrI) and the PDE (thin solid curve = median, dashed curves = 90% PrI). Vertical dashed lines in middle column show the initial location of tagged agents. Discrete parameter values $M=1$ and $\rho_{x}=Q=0$ corresponding to $D=0.25$ and $v=\lambda=0$ for a simulation with $\Delta=\tau=1$.
3.3 Unbiased motility and proliferation
Including proliferation in the ABM, with no directional bias, means that the total population size increases over time and the mean-field PDE model is the well-known Fisher-Kolmogorov model [36, 37, 38]. Therefore, after a sufficient duration of time the macroscopic agent density eventually approaches a travelling wave profile with $C(x,t)\to 1^{-}$ (representing a fully occupied lattice) behind the wave front and $C(x,t)\to 0^{+}$ ahead of the leading edge (Figure 2g). The distributions of tagged agents initially close to either leading edge start to spread out at the beginning of the simulation, and eventually become skewed away from the centre (Figure 2h, blue and yellow), as occurred in the case without proliferation in Figure 2a-c. However, these agents that are initially at the leading-ledge of the population eventually become relatively immobile and fixed in place once the lattice in their local neighbourhood becomes fully occupied. As a result, there is little further change in their distributions after approximately $t=150$ (Figure 2i). The tagged agents initially located in the centre of the population near $x=0$ are unable to move at all and so retain their initial Dirac delta-like distribution for the whole simulation (Figure 2h-i, red). Again, the solution of the PDE model for $P(x,t)$ captures these trends and dynamics reasonably accurately.
3.4 Biased motility and proliferation
In simulations with both directional bias and proliferation, the macroscopic agent density develops an asymmetric profile, with the travelling wave at the right-most leading edge moving faster than that at the left-most leading edge (Figure 2j). Tagged agent distributions display varying degrees of drift and diffusion depending on their initial location within the population (Figure 2k). Tagged agents in the centre are unable to move and are fixed in place almost immediately due to high local density (Figure 2k-l, red). Those tagged agents initially near the left-hand edge spread out until around $t=150$ when they become immobile (Figure 2k-l, blue), whereas the tagged agents initially near the right-most leading edge are able to move more rapidly and their distribution is continuing to spread out at the end of the simulation at $t=300$ (Figure 2k-l, yellow). This is because some of these tagged agents are able to remain ahead of the wave front in a region where macroscopic density remains sufficiently small. Again the PDE provides a good approximation to the ABM results.
3.5 Accuracy of the approximation for tagged agents
We explored how the accuracy of the PDE model compared to the method proposed by Simpson, Landman and Hughes [27] for calculating the mean and variance of tagged agent locations. This method, which we refer to as the SLH approximation, relies on calculating the rate of change of the mean and variance for an agent that is positioned at the mean location, and thus ignores the effects of uncertainty in the agent location. In all cases investigated, the PDE model and the SLH approximation both predict the mean agent location very accurately (Supplementary Figures S1a,c,e,g [34]). In cases without proliferation, the PDE model predicts the variance as well or better than the SLH approximation (Supplementary Figure S1b,d [34]). When there is proliferation, the results are more mixed. When there is no bias, the SLH predicts the variance of agents initially near a leading edge more accurately than the PDE model (Supplementary Figure S1f [34]), but the reverse is true when there is bias (Supplementary Figure S1h [34]). The PDE model tends to overestimate the variance in tagged locations in all cases, while the SLH approximation underestimates it in some cases and overestimates in others.
Whilst we have calculated the SLH approximation for all cases investigated here for completeness, it should be noted that Simpson et al. [27] did not consider proliferation, and only examined agents initially located near the right-most leading edge of the population. Also the SLH approximation provides no information about tagged agent locations beyond their mean and variance, whereas our model provides access to the full distribution via its PDF. This provides a richer set of information, which enables predictions and analyses that would not otherwise be possible. For example, our results could be used to define a tractable likelihood function for observed data on tagged agent locations. This would enable parameter inference and uncertainty quantification via either frequentist or Bayesian methods [33].
To illustrate this, we explored a test case in which all lattice sites were initially occupied with probability $0.5$ and there is no proliferation or motility bias. This means that the macroscopic agent density is uniformly constant at $C(x,t)=0.5$ and the PDE for $P(x,t)$ in Equation (8) reduces to the linear diffusion equation with diffusivity $0.5D$. This PDE for $P(x,t)$ would accurately describe a situation where 50% of the tagged agent’s attempted movement events are aborted. However, simulations of the ABM show that the tagged agent’s distribution spreads out more slowly than this PDE predicts (Supplementary Figure S2 [34]), indicating that more than 50% of moves are aborted. This is due to the positive correlation between neighbouring lattice sites, which means that, conditional on the tagged agent being at site $(i,j)$, the occupancy probabilities of sites $(i\pm 1,j)$ and $(i,j\pm 1)$ are slightly greater than $0.5$.
4 Conclusion and Future Work
In this work, we have investigated lattice-based random walk models of individual-level motility and proliferation mechanisms in the context of an exclusion process framework, where each lattice site can be occupied by no more than a single agent [21]. Stochastic models with exclusion are often used to represent cell biology experiments where crowding effects can be very pronounced [1, 2, 3, 4, 7, 8, 9, 10, 11, 13, 14, 15]. Motivated by experimental observations in Figure 1, we used a stochastic ABM to describe the spatial evolution of the population-level density profile, as well as considering the motion of a small number of tagged agents within the broader population. We explored the existing continuum-limit description of the macroscopic density $C(x,y,t)\in[0,1]$ alongside a new continuum-limit description of the PDF for the location of tagged agents $P(x,y,t)\geq 0$. The derivation of the continuum limit was extended to the situation where we have $S\in\mathbb{N}^{+}$ distinct subpopulations on the lattice so that the dependent variables in the PDE models are $C^{(s)}(x,y,t)\in[0,1]$ and $P^{(s)}(x,y,t)\geq 0$ for $s=1,2,3,\ldots,S$. Repeated stochastic simulation data showed that the continuum-limit model for $C$ and $P$ provide a good match data obtained from the computationally expensive stochastic simulations. In particular, we showed that numerical estimates of $P$ provide a probabilistic interpretation of the motion of tagged agents since the continuum-limit model can be used to predict both the expected location of the tagged agents, as well as predicting the variability in their location. Numerical tests confirmed that the solution of the continuum-limit model for $P$ provides a reasonable match to data from the stochastic ABM under a range of conditions including unbiased and biased motility, both in the presence and absence of agent proliferation.
The PDE models we derived for the density $C^{(s)}$ involve a source term whenever $Q^{(s)}>0$ for $s=1,2,\ldots,S$. The corresponding PDE for the PDF for tagged agents $P^{(s)}$ does not involve any source term, even when the simulation involves agent proliferation with $Q^{(s)}>0$. This property is a reflection of the fact that our model tracks the location of agents that are tagged at the beginning of the experiment, without consideration of any offspring they may produce. Our approach could be generalised to track the original tagged agents, as well as any daughter agents and associated lineages they give rise to. This would lead to a PDE model for $P^{(s)}$ that involves a source term to account for the fact that tagged agents themselves can be involved in proliferation events, similar to mathematical models of lineage tracing [39].
A potential use case for our framework is particle identification. For example, suppose there are multiple tagged agents that are are observed at snapshots in time, but are not a priori distinguishable. Our model provides the probability that an agent which was initially at known location ${\bf x}_{0}$ will be located at location $\bf x$} at time $t$. This could be used to provide a probabilistic estimate of an observed agent’s identity, but we leave this question for future work. Another approach for generalising our mathematical models is to recast the stochastic lattice-based random walk model into a stochastic lattice-free random walk [40, 41, 42]. The main advantage of working in a lattice-free framework is that the motility direction is continuous rather than being discrete, however this additional flexibility comes at the cost of additional mathematical complexity, which makes deriving appropriate mean-field PDE descriptions more difficult [40, 41, 42].
Regardless of whether we work with a lattice-based or lattice-free framework, deriving appropriate continuum limits for the population-level density and individual-level tagged agent trajectory properties provides an opportunity to perform parameter inference using simultaneous observations of the agent density, $C^{(s)}$, and the location of tagged agents, $P^{(s)}$ for $s=1,2,\ldots,S$. Having simple computational tools that enable us to understand how $C^{(s)}$ and $P^{(s)}$ vary with the model parameters provides a key ingredient for either Bayesian or frequentist statistical inference [43, 44, 45] that makes full use of all available (combined) data. Having a continuum-limit PDE description is highly advantageous for parameter inference as: (1) when coupled with a suitable observation noise model [33, 43], it provides access to a likelihood function that can be used for optimisation or sampling-based methods; and (2) it is more computationally efficient than generating repeated stochastic ABM simulations, which can be prohibitively expensive for performing parameter inference, especially if the parameter space is high-dimensional.
The approach we have taken in this study takes the intrinsic movement, proliferation and interaction rules that define the ABM as given and uses these to make predictions about how population density and agent locations evolve over time. In some situations, these rules may be approximately known from biophysical principles or prior data. In other situations, they may be unknown, in which case models such as those presented here can be used to statistically test alternative hypotheses against observed data, using parameter inference and model selection approaches [43, 33, 46]. Hence, our methods provide an avenue to inferring mechanistic behavioural rules from empirical cell trajectory data.
Acknowledgements This work is partly supported by the Australian Research Council (DP230100025, CE230100001) and the Marsden Fund (24-UOC-020). The authors are grateful to two anonymous reviewers for comments on an earlier version of this article.
Supplementary Material
S1 Numerical methods